Please help me to sketch the following line by method of Linear Function or straight lines.
7. y = (2/3)x-4
8. 8x-3y=24
9. slope=-3/4
Thanks, I am grateful for your help.

Answer by checkley71(8403)

Y=(2/3)X-4 THIS REPRESENTS A LINE WITH A SLOPE OF 2/3 & A Y INTERCEPT OF (0,-4)
TO FIND A SECOND POINT ON THIS LINE SET Y=0 & SOLVE FOR X THUS
0=(2/3)X-4
(2/3)X=4
X=4*3/2
X=6 OR (6,0) IS A SECOND POINT.
---------------------------------------------------------------------------
8X-3Y=24
-3Y=-8X+24
Y=(-8/-3)X+24/-3
Y=(8/3)X-8 THIS IS A LINE WITH A SLOPE OF 8/3 & A Y INTERCEPT OF -8 (0,-8)
AGAIN TO FIND THE X INTERCEPT SET Y=0
0=(8/3)X-8
(8/3)X=8
X=8*3/8
X=3
----------------------------------------------------
SLOPE=-3/4 IS A LINE THAT GOES FROM A POINT TO THE LEFT (NEGATIVE SLOPE) & UP.
FOR EVERY 4 UNITS LEFT YOU NEED TO GO 3 UNITS UP.